# Code Readability in the Age of Large Language Models: An Industrial Case Study from Atlassian

### v2

@@ 4.1. Approach
The empirical study aims to understand code readability via static code analysis. In this section, we describe the design of our case study on HULA (takerngsaksiri2024human) over our internal dataset.
@@
Figure 3. The Overview of the RQ2 Empirical Study Comparing Human-written Code to HULA-generated Code on Code Readability Measures.
@@
Overview. Figure 3 shows the overview of the approach in this research question. Given a JIRA issue and a corresponding code repository, a software engineer writes a code, while AI coding agents, HULA, generate the codes to resolve the JIRA issue. Then, the human-written code and the HULA-generated code are assessed via eight code readability measurements. Lastly, we compare the results using the statistical analysis of the significance of Mann-Whitney U Test (mann1947test) and the effect size of Cliff’s Delta (cliff1993dominance).
@@
Large Language Model. We evaluate the code readability using our Human-in-the-loop Software Development Agents framework, HULA (takerngsaksiri2024human). HULA is Atlassian’s LLM-based agents framework for software development available internally in Jira Software. Specifically, the framework is used to generate a code patch given a code repository and a description of the Jira issue to be resolved. In this version, HULA uses GPT-4 (openai2023gpt4) as a based LLM and the framework is evaluated without additional human feedback.
@@
Datasets. We evaluate HULA via the internal dataset (takerngsaksiri2024human), which includes a set of completed software development tasks (i.e., Jira issues, repositories and corresponding pull requests from BitBucket) at Atlassian. The dataset consists of 144 Jira issues with 250 coding files spanning six programming languages: TypeScript, Java, Kotlin, Python, Go and Scala. The distribution of programming languages is illustrated in Figure 4.
Figure 4. The Distribution of Programming Languages in the Dataset (n=250).
@@
Code Readability Measurement. Following the recent study on code readability (al2022readable), we cautiously select the following eight static analysis metrics that are language-agnostic (i.e., can be used in multilingual) via the Multimetric library(multimetric2024).
@@
First, we use Line of Code to evaluate the total number of lines of code in the file, expressing the file size. The Comment Ratio is used to evaluate the percentage of the lines of comment to code, expressing the descriptive of the code. Then, Cyclomatic Complexity (mccabe1976complexity) is used to evaluate the number of decision paths of the code, expressing the code complexity. Next, Maintainability Index (Oman1992maintainability) is used to measure how maintainable the code is. Lastly, we utilize Halstead Metrics (halstead1977elements), a set of metrics designed to quantify various aspects of code. These metrics are employed to evaluate key attributes, including Difficulty (the difficulty of the program to write or understand), Vocabulary (the number of unique operators and unique operands occurrence), Volume (the program length and vocabulary size), and Time Required (the estimated time for implementation).
@@
Figure 5. (RQ2) The Scores of Code Readability Measures between HULA-generated and Human-written Code.
Figure 6. (RQ2) The Median Scores of Code Readability Measures between Human-written and HULA-generated code by Programming Languages.
  1. (1)
@@
  3. (3)
    Cyclomatic Complexity (mccabe1976complexity) or McCabe number is the number of decisions path of code according. Specifically, the complexity is defined as $C=E-N+2$, where $E$ is the number of edges on the control flow graph and $N$ is the number of nodes on the control flow graph.
  4. (4)
    Maintainability Index (Oman1992maintainability) is a metric to measure how maintainable the code is. We use the Microsoft formula (microsoft2022maintainability) defined as $Max(0,(171-5.2*ln(HalsteadVolume)-0.23*(McCabe)-16.2*ln(LoC))*100/171)$.
  5. (5)
    Halstead Difficulty (halstead1977elements) defines as $D=(n1/2)*(N2/n2)$, where $n1$ is the number of distinct operators, $n2$ is the number of distinct operands, $N1$ is the total number of operators, and $N2$ is the total number of operands.
  6. (6)
    Halstead Vocabulary (halstead1977elements) defines as $n=n1+n2$.
  7. (7)
    Halstead Volume (halstead1977elements) defines as $V=(N1+N2)*log_{2}(n)$.
  8. (8)
    Halstead Time Required (halstead1977elements) defines as $T=(D*V)/18$
4.2. Results
Figure 5 shows the results of this research question, covering the case study of HULA on our internal dataset. Table 8 shows the statistical significance and the effect size of the result. Lastly, Figure 6 shows the detailed analysis of code readability measurement by programming languages.
The generated code is comparable to the human-written code with a negligible to small statistical difference. Figure 5 shows that HULA, with a GPT-4 based model, generates code that is highly similar to humans with only a slightly longer code but with no statistically significant difference in Cyclomatic Complexity and comment ratio. On Halstead’s metrics, our LLM-generated code has marginally higher values on difficulty, vocabulary, volume, and time required. These results indicate that the HULA framework can produce code that closely resembles human-written code in terms of code readability, indicating the promising performance of LLM in production.
@@ 4.2. Results
Table 8. (RQ2) The mean significance from Mann-Whitney U Test (mann1947test) and Effect Size measured by the Cliff’s Delta (cliff1993dominance) based on the comparison of HULA and Human code.
@@ 5. Threat to Validity
Threat to construct validity. We use an empirical study with static code analysis as a proxy for code readability measurement between LLM-generated and Human code. We cautiously follow the existing work on selecting the metrics representing code readability. However, we acknowledge the limitations that the metrics might not completely represent the code readability from all perspectives.

### v3

@@ V-A Approach
The empirical study aims to understand code readability via static code analysis. In this section, we describe the design of our case study on HULA [23] over our internal dataset.
@@
Figure 4: The Overview of the RQ2 Empirical Study Comparing Human-written Code to HULA-generated Code on Code Readability Measures.
@@
Overview. Figure 4 shows the overview of the approach in this research question. Given a JIRA issue and a corresponding code repository, a software engineer writes a code, while AI coding agents, HULA, generate the codes to resolve the JIRA issue. Then, the human-written code and the HULA-generated code are assessed via eight code readability measurements. Lastly, we compare the results using the statistical analysis of the significance of Mann-Whitney U Test [50] and the effect size of Cliff’s Delta [51].
@@
Large Language Model. We evaluate the code readability using our Human-in-the-loop Software Development Agents framework, HULA [23]. HULA is Atlassian’s LLM-based agents framework for software development available internally in Jira Software. Specifically, the framework is used to generate a code patch given a code repository and a description of the Jira issue to be resolved. In this version, HULA uses GPT-4 [24] as a based LLM and the framework is evaluated without additional human feedback.
@@
Datasets. We evaluate HULA via the internal dataset [23], which includes a set of completed software development tasks (i.e., Jira issues, repositories and corresponding pull requests from BitBucket) at Atlassian. The dataset consists of 144 Jira issues with 250 coding files spanning six programming languages: TypeScript, Java, Kotlin, Python, Go and Scala. The distribution of programming languages is illustrated in Figure 5.
Figure 5: The Distribution of Programming Languages in the Dataset (n=250).
@@
Code Readability Measurement. Following the recent study on code readability [40], we cautiously select the following eight static analysis metrics that are language-agnostic (i.e., can be used in multilingual) via the Multimetric library[52].
@@
First, we use Line of Code to evaluate the total number of lines of code in the file, expressing the file size. The Comment Ratio is used to evaluate the percentage of the lines of comment to code, expressing the descriptive of the code. Then, Cyclomatic Complexity [53] is used to evaluate the number of decision paths of the code, expressing the code complexity. Next, Maintainability Index [54] is used to measure how maintainable the code is. Lastly, we utilize Halstead Metrics [55], a set of metrics designed to quantify various aspects of code. These metrics are employed to evaluate key attributes, including Difficulty (the difficulty of the program to write or understand), Vocabulary (the number of unique operators and unique operands occurrence), Volume (the program length and vocabulary size), and Time Required (the estimated time for implementation).
@@
Figure 6: (RQ2) The Scores of Code Readability Measures between HULA-generated and Human-written Code.
Figure 7: (RQ2) The Median Scores of Code Readability Measures between Human-written and HULA-generated code by Programming Languages.
  1. 1.
@@
  3. 3.
    Cyclomatic Complexity [53] or McCabe number is the number of decisions path of code according. Specifically, the complexity is defined as $C=E-N+2$, where $E$ is the number of edges on the control flow graph and $N$ is the number of nodes on the control flow graph.
  4. 4.
    Maintainability Index [54] is a metric to measure how maintainable the code is. We use the Microsoft formula [56] defined as $Max(0,(171-5.2*ln(HalsteadVolume)-0.23*(McCabe)-16.2*ln(LoC))*100/171)$.
  5. 5.
    Halstead Difficulty [55] defines as $D=(n1/2)*(N2/n2)$, where $n1$ is the number of distinct operators, $n2$ is the number of distinct operands, $N1$ is the total number of operators, and $N2$ is the total number of operands.
  6. 6.
    Halstead Vocabulary [55] defines as $n=n1+n2$.
  7. 7.
    Halstead Volume [55] defines as $V=(N1+N2)*log_{2}(n)$.
  8. 8.
    Halstead Time Required [55] defines as $T=(D*V)/18$
V-B Results
Figure 6 shows the results of this research question, covering the case study of HULA on our internal dataset. Table VIII shows the statistical significance and the effect size of the result. Lastly, Figure 7 shows the detailed analysis of code readability measurement by programming languages.
The HULA-generated code is comparable to the human-written code with a negligible to small statistical difference across our eight evaluation metrics. Figure 6 shows that HULA, with a GPT-4 based model, generates code that is highly similar to humans with only a slightly longer code but with no statistically significant difference in Cyclomatic Complexity and comment ratio. On Halstead’s metrics, our LLM-generated code has marginally higher values on difficulty, vocabulary, volume, and time required. These results indicate that the HULA framework can produce code that closely resembles human-written code in terms of code readability, indicating the promising performance of LLM in production.
@@ V-B Results
TABLE VIII: (RQ2) The mean significance from Mann-Whitney U Test [50] and Effect Size measured by the Cliff’s Delta [51] based on the comparison of HULA and Human code.
@@ VI Threat to Validity
Threat to construct validity. We use an empirical study with static code analysis as a proxy for code readability measurement between LLM-generated and Human code. We cautiously follow the existing work on selecting the metrics representing code readability. However, we acknowledge the limitations that the metrics might not completely represent the code readability from all perspectives and that some inter-dependencies may exist among the metrics.
@@ VII Conclusion and Implications
VIII Disclaimer
The perspectives and conclusions presented in this paper are solely the authors’ and should not be interpreted as representing the official policies or endorsements of Atlassian or any of its subsidiaries and affiliates. Additionally, the outcomes of this paper are independent of, and should not be construed as an assessment of, the quality of products offered by Atlassian.